# Graph Retrieval-Augmented Generation: A Survey

### v1

@@ Abstract.
Recently, Retrieval-Augmented Generation (RAG) has achieved remarkable success in addressing the challenges of Large Language Models (LLMs) without necessitating retraining. By referencing an external knowledge base, RAG refines LLM outputs, effectively mitigating issues such as “hallucination”, lack of domain-specific knowledge, and outdated information. However, the complex structure of relationships among different entities in databases presents challenges for RAG systems. In response, GraphRAG leverages structural information across entities to enable more precise and comprehensive retrieval, capturing relational knowledge and facilitating more accurate, context-aware responses. Given the novelty and potential of GraphRAG, a systematic review of current technologies is imperative. This paper provides the first comprehensive overview of GraphRAG methodologies. We formalize the GraphRAG workflow, encompassing Graph-Based Indexing, Graph-Guided Retrieval, and Graph-Enhanced Generation. We then outline the core technologies and training methods at each stage. Additionally, we examine downstream tasks, application domains, evaluation methodologies, and industrial use cases of GraphRAG. Finally, we explore future research directions to inspire further inquiries and advance progress in the field.
@@
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Computing methodologies Knowledge representation and reasoningccs: Information systems Information retrievalccs: Information systems Data mining
@@ 1. Introduction
The development of Large Language Models like GPT-4 (OpenAI, 2024), Qwen2 (Yang et al., 2024c), and LLaMA (Dubey et al., 2024) has sparked a revolution in the field of artificial intelligence, fundamentally altering the landscape of natural language processing. These models, built on Transformer (Vaswani et al., 2017) architectures and trained on diverse and extensive datasets, have demonstrated unprecedented capabilities in understanding, interpreting, and generating human language. The impact of these advancements is profound, stretching across various sectors including healthcare (Wang et al., 2024c; Liu et al., 2024b; Zheng et al., 2024), finance (Nie et al., 2024; Li et al., 2024d), and education (Wang et al., 2024d; Ghimire et al., 2024), where they facilitate more nuanced and efficient interactions between humans and machines.
Despite their remarkable language comprehension and text generation capabilities, LLMs may exhibit limitations due to a lack of domain-specific knowledge, real-time updated information, and proprietary knowledge, which are outside LLMs’ pre-training corpus. These gaps can lead to a phenomenon known as “hallucination” (Huang et al., 2023b) where the model generates inaccurate or even fabricated information. Consequently, it is imperative to supplement LLMs with external knowledge to mitigate this problem. Retrieval-Augmented Generation (RAG) (Fan et al., 2024a; Hu and Lu, 2024; Huang and Huang, 2024; Zhao et al., 2024; Gao et al., 2024b; Yu et al., 2024; Wu et al., 2024a) emerged as a significant evolution, which aims to enhance the quality and relevance of generated content by integrating a retrieval component within the generation process. The essence of RAG lies in its ability to dynamically query a large text corpus to incorporate relevant factual knowledge into the responses generated by the underlying language models. This integration not only enriches the contextual depth of the responses but also ensures a higher degree of factual accuracy and specificity. RAG has gained widespread attention due to its exceptional performance and broad applications, becoming a key focus within the field.
@@ 2.1. RAG
RAG combines external knowledge with LLMs for improved task performance, integrating domain-specific information to ensure factuality and credibility. In the past two years, researchers have written many comprehensive surveys about RAG (Fan et al., 2024a; Hu and Lu, 2024; Huang and Huang, 2024; Zhao et al., 2024; Gao et al., 2024b; Yu et al., 2024; Wu et al., 2024a). For example, Fan et al. (2024a) and Gao et al. (2024b) categorize RAG methods from the perspectives of retrieval, generation, and augmentation. Zhao et al. (2024) review RAG methods for databases with different modalities. Yu et al. (2024) systematically summarize the evaluation of RAG methods. These works provide a structured synthesis of current RAG methodologies, fostering a deeper understanding and suggesting future directions of the area.
@@ 2.2. LLMs on Graphs
LLMs are revolutionizing natural language processing due to their excellent text understanding, reasoning, and generation capabilities, along with their generalization and zero-shot transfer abilities. Although LLMs are primarily designed to process pure text and struggle with non-Euclidean data containing complex structural information, such as graphs (Wang et al., 2023b; Guo et al., 2023), numerous studies (Fan et al., 2024b; Mao et al., 2024c; Liu et al., 2024c; Li et al., 2024c; Pan et al., 2023, 2024; Jin et al., 2024a; Chen, 2024; Zhu et al., 2024; Wang et al., 2024e) have been conducted in these fields. These papers primarily integrate LLMs with GNNs to enhance modeling capabilities for graph data, thereby improving performance on downstream tasks such as node classification, edge prediction, graph classification, and others. For example, Zhu et al. (2024) propose an efficient fine-tuning method named ENGINE, which combines LLMs and GNNs through a side structure for enhancing graph representation.
@@ 2.3. KBQA
KBQA is a significant task in natural language processing, aiming to respond to user queries based on external knowledge bases (Fu et al., 2020; Lan et al., 2023, 2021; Yani and Krisnadhi, 2021), thereby achieving goals such as fact verification, passage retrieval enhancement, and text understanding. Previous surveys typically categorize existing KBQA approaches into two main types: Information Retrieval (IR)-based methods and Semantic Parsing (SP)-based methods. Specifically, IR-based methods (Luo et al., 2024b; Sun et al., 2024b; Jiang et al., 2023b; Zhang et al., 2022b; Wu et al., 2023b; Wang et al., 2023a; Jiang et al., 2024b) retrieve information related to the query from the knowledge graph (KG) and use it to enhance the generation process. While SP-based methods (Chakraborty, 2024; Fang et al., 2024; Chen et al., 2021; Ye et al., 2021; Gu and Su, 2022; Sun et al., 2023) generate a logical form (LF) for each query and execute it against knowledge bases to obtain the answer.
@@ 5.2.1. Graph Indexing
Graph indexing represents the most commonly used approach, preserving the entire structure of the graph. This method ensures that for any given node, all its edges and neighboring nodes are easily accessible. During subsequent retrieval stages, classic graph search algorithms such as BFS and Shortest Path Algorithms can be employed to facilitate retrieval tasks (Jin et al., 2024b; Luo et al., 2024b; Sun et al., 2024b; Jiang et al., 2024a; Yasunaga et al., 2021; Taunk et al., 2023).
@@ 5.2.3. Vector Indexing
Remark. These three indexing methods each offer distinct advantages: graph indexing facilitates easy access to structural information, text indexing simplifies retrieval of textual content, and vector indexing enables quick and efficient searches. Therefore, in practical applications, a hybrid approach combining these indexing methods is often preferred over relying solely on one.
@@ 6.1.3. GNN-based Retriever
GNNs are adept at understanding and leveraging complex graph structures. GNN-based retrievers typically encode graph data and subsequently score different retrieval granularities based on their similarity to the query. For example, GNN-RAG (Mavromatis and Karypis, 2024) first encodes the graph, assigns a score to each entity, and retrieves entities relevant to the query based on a threshold. EtD (Liu et al., 2024d) iterates multiple times to retrieve relevant paths. During each iteration, it first uses LLaMA2 (Touvron et al., 2023) to select edges connecting the current node, then employs GNNs to obtain embeddings of the new layer of nodes for the next round of LLM selection.
Remark. During the retrieval process, non-parametric retrievers exhibit good retrieval efficiency, but they may suffer from inaccurate retrieval due to a lack of training on downstream tasks. Meanwhile, although LM-based retrievers and GNN-based retrievers offer higher retrieval accuracy, they require significant computational overhead. Considering this complementarity, many methods propose hybrid retrieval approaches to improve both retrieval efficiency and accuracy. Many approaches adopt a multi-stage retrieval strategy, employing different models at each stage. For example, RoG (Luo et al., 2024b) first utilizes LLMs to generate planning paths and then extracts paths satisfying the planning paths from knowledge graphs. GenTKGQA (Gao et al., 2024a) infers crucial relations and constraints from the query using LLMs and extracts triplets according to these constraints.
@@ 6.2.2. Iterative Retrieval
One distinctive characteristic of adaptive retrieval is to let models autonomously determine the optimal moments to finish the retrieval activities. For instance, (Wu et al., 2023b; Guo et al., 2024) leverage an LM for hop prediction, which serves as an indicator to end the retrieval. There is also a group of researchers who utilize model-generated special tokens or texts as termination signals for the retrieval process. For example, ToG (Sun et al., 2024b) prompts the LLM agent to explore the multiple possible reasoning paths until the LLM determines the question can be answered based on the current reasoning path. (Zhang et al., 2022b) trains a RoBERTa to expand a path from each topic entity. In the process, a virtual relation named as “[END]” is introduced to terminate the retrieval process.
@@ 6.2.3. Multi-Stage Retrieval
Remark. In GraphRAG, once retrieval typically exhibits lower complexity and shorter response times, making it suitable for scenarios requiring real-time responsiveness. In contrast, iterative retrieval often involves higher time complexity, especially when employing LLMs as retrievers, potentially leading to longer processing times. However, this approach can yield higher retrieval accuracy by iteratively refining retrieved information and generating responses. Therefore, the choice of retrieval paradigm should balance accuracy and time complexity based on specific use cases and requirements.
@@ 6.3.1. Nodes
Nodes allow for precise retrieval focused on individual elements within the graph, which is ideal for targeted queries and specific information extraction. In general, for knowledge graphs, nodes refer to entities. For other types of text attribute graphs, nodes may include textual information that describes the node’s attributes. By retrieving nodes within the graph, GraphRAG systems could provide detailed insights into their attributes, relationships, and contextual information. For example, Munikoti et al. (2023), Li et al. (2024b) and Wang et al. (2024b) construct document graphs and retrieves relevant passage nodes. Liu et al. (2024d), Sun et al. (2019) and Gutiérrez et al. (2024) retrieve entities from constructed knowledge graphs.
@@ 6.3.3. Paths
However, path retrieval can be challenging due to the exponential growth in possible paths as graph size increases, which escalates computational complexity. To address this, some methods retrieve relevant paths based on pre-defined rules. For example, Wang et al. (2024a) and Lo and Lim (2023) first select entity pairs in the query and then traverse to find all the paths between them within $n$-hop. HyKGE (Jiang et al., 2024a) first defines three types of paths: path, co-ancestor chain, and co-occurrence chain, and then utilizes corresponding rules to retrieve each of these three types of paths. In addition, some methods utilize models to perform path searching on graphs. ToG (Sun et al., 2024b) proposes to prompt the LLM agent to perform the beam search on KGs and find multiple possible reasoning paths that help answer the question. Luo et al. (2024b), Wu et al. (2023b) and Guo et al. (2024) first utilizes the model to generate faithful reasoning plans and then retrieves relevant paths based on these plans. GNN-RAG (Mavromatis and Karypis, 2024) first identifies the entities in the question. Subsequently, all paths between entities that satisfy a certain length relationship are extracted.
@@ 6.3.5. Hybrid Granularties
Remark. (1) In real applications, there are no clear boundaries between these retrieval granularities, as subgraphs can be composed of multiple paths, and paths can be formed by several triplets. (2) Various granularities such as nodes, triplets, paths, and subgraphs offer distinct advantages in the GraphRAG process. Balancing between retrieval content and efficiency is crucial when selecting the granularity, depending on the specific context of the task. For straightforward queries or when efficiency is paramount, finer granularities such as entities or triplets may be preferred to optimize retrieval speed and relevance. In contrast, complex scenarios often benefit from a hybrid approach that combines multiple granularities. This approach ensures a more comprehensive understanding of the graph structure and relationships, enhancing the depth and accuracy of the generated responses. Thus, GraphRAG’s flexibility in granularity selection allows it to adapt effectively to diverse information retrieval needs across various domains.
@@ 6.4.1. Query Enhancement
Due to the generally short length of queries and their limited information content, query expansion aims to improve search results by supplementing or refining the original query with additional relevant terms or concepts. Luo et al. (2024b) generate relation paths grounded by KGs with LLMs to enhance the retrieval query. Cheng et al. (2024) adopt SPARQL to get all the aliases of the query entities from Wikidata to augment the retrieval queries, which capture lexical variations of the same entity. Huang et al. (2023a) propose a consensus-view knowledge retrieval method to improve retrieval accuracy, which first discover semantically relevant queries, and then re-weight the original query terms to enhance the retrieval performance. HyKGE (Jiang et al., 2024a) utilizes a large model to generate the hypothesis output of the question, concatenating the hypothesis output with the query as input to the retriever.

### v2

@@ Abstract.
Recently, Retrieval-Augmented Generation (RAG) has achieved remarkable success in addressing the challenges of Large Language Models (LLMs) without necessitating retraining. By referencing an external knowledge base, RAG refines LLM outputs, effectively mitigating issues such as “hallucination”, lack of domain-specific knowledge, and outdated information. However, the complex structure of relationships among different entities in databases presents challenges for RAG systems. In response, GraphRAG leverages structural information across entities to enable more precise and comprehensive retrieval, capturing relational knowledge and facilitating more accurate, context-aware responses. Given the novelty and potential of GraphRAG, a systematic review of current technologies is imperative. This paper provides the first comprehensive overview of GraphRAG methodologies. We formalize the GraphRAG workflow, encompassing Graph-Based Indexing, Graph-Guided Retrieval, and Graph-Enhanced Generation. We then outline the core technologies and training methods at each stage. Additionally, we examine downstream tasks, application domains, evaluation methodologies, and industrial use cases of GraphRAG. Finally, we explore future research directions to inspire further inquiries and advance progress in the field. In order to track recent progress in this field, we set up a repository at https://github.com/pengboci/GraphRAG-Survey.
@@
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 9ccs: Computing methodologies Knowledge representation and reasoningccs: Information systems Information retrievalccs: Information systems Data mining
@@ 1. Introduction
The development of Large Language Models like GPT-4 (OpenAI, 2024), Qwen2 (Yang et al., 2024c), and LLaMA (Dubey et al., 2024) has sparked a revolution in the field of artificial intelligence, fundamentally altering the landscape of natural language processing. These models, built on Transformer (Vaswani et al., 2017) architectures and trained on diverse and extensive datasets, have demonstrated unprecedented capabilities in understanding, interpreting, and generating human language. The impact of these advancements is profound, stretching across various sectors including healthcare (Wang et al., 2024c; Liu et al., 2024c; Zheng et al., 2024), finance (Nie et al., 2024; Li et al., 2024d), and education (Wang et al., 2024d; Ghimire et al., 2024), where they facilitate more nuanced and efficient interactions between humans and machines.
Despite their remarkable language comprehension and text generation capabilities, LLMs may exhibit limitations due to a lack of domain-specific knowledge, real-time updated information, and proprietary knowledge, which are outside LLMs’ pre-training corpus. These gaps can lead to a phenomenon known as “hallucination” (Huang et al., 2023b) where the model generates inaccurate or even fabricated information. Consequently, it is imperative to supplement LLMs with external knowledge to mitigate this problem. Retrieval-Augmented Generation (RAG) (Fan et al., 2024a; Hu and Lu, 2024; Huang and Huang, 2024; Zhao et al., 2024; Gao et al., 2024b; Yu et al., 2024; Wu et al., 2024b) emerged as a significant evolution, which aims to enhance the quality and relevance of generated content by integrating a retrieval component within the generation process. The essence of RAG lies in its ability to dynamically query a large text corpus to incorporate relevant factual knowledge into the responses generated by the underlying language models. This integration not only enriches the contextual depth of the responses but also ensures a higher degree of factual accuracy and specificity. RAG has gained widespread attention due to its exceptional performance and broad applications, becoming a key focus within the field.
@@ 2.1. RAG
RAG combines external knowledge with LLMs for improved task performance, integrating domain-specific information to ensure factuality and credibility. In the past two years, researchers have written many comprehensive surveys about RAG (Fan et al., 2024a; Hu and Lu, 2024; Huang and Huang, 2024; Zhao et al., 2024; Gao et al., 2024b; Yu et al., 2024; Wu et al., 2024b). For example,  Fan et al. (2024a) and Gao et al. (2024b) categorize RAG methods from the perspectives of retrieval, generation, and augmentation.  Zhao et al. (2024) review RAG methods for databases with different modalities.  Yu et al. (2024) systematically summarize the evaluation of RAG methods. These works provide a structured synthesis of current RAG methodologies, fostering a deeper understanding and suggesting future directions of the area.
@@ 2.2. LLMs on Graphs
LLMs are revolutionizing natural language processing due to their excellent text understanding, reasoning, and generation capabilities, along with their generalization and zero-shot transfer abilities. Although LLMs are primarily designed to process pure text and struggle with non-Euclidean data containing complex structural information, such as graphs (Wang et al., 2023b; Guo et al., 2023), numerous studies (Fan et al., 2024b; Mao et al., 2024c; Liu et al., 2024d; Li et al., 2024c; Pan et al., 2023, 2024; Jin et al., 2024a; Chen, 2024; Zhu et al., 2024; Wang et al., 2024e) have been conducted in these fields. These papers primarily integrate LLMs with GNNs to enhance modeling capabilities for graph data, thereby improving performance on downstream tasks such as node classification, edge prediction, graph classification, and others. For example,  Zhu et al. (2024) propose an efficient fine-tuning method named ENGINE, which combines LLMs and GNNs through a side structure for enhancing graph representation.
@@ 2.3. KBQA
KBQA is a significant task in natural language processing, aiming to respond to user queries based on external knowledge bases (Fu et al., 2020; Lan et al., 2023, 2021; Yani and Krisnadhi, 2021), thereby achieving goals such as fact verification, passage retrieval enhancement, and text understanding. Previous surveys typically categorize existing KBQA approaches into two main types: Information Retrieval (IR)-based methods and Semantic Parsing (SP)-based methods. Specifically, IR-based methods (Luo et al., 2024b; Sun et al., 2024b; Ma et al., 2024; Jiang et al., 2023b; Zhang et al., 2022b; Wu et al., 2023b; Wang et al., 2023a; Jiang et al., 2024b) retrieve information related to the query from the knowledge graph (KG) and use it to enhance the generation process. While SP-based methods (Chakraborty, 2024; Fang et al., 2024b; Chen et al., 2021; Ye et al., 2021; Gu and Su, 2022; Sun et al., 2023) generate a logical form (LF) for each query and execute it against knowledge bases to obtain the answer.
@@ 5.2.1. Graph Indexing
Graph indexing represents the most commonly used approach, preserving the entire structure of the graph. This method ensures that for any given node, all its edges and neighboring nodes are easily accessible. During subsequent retrieval stages, classic graph search algorithms such as BFS and Shortest Path Algorithms can be employed to facilitate retrieval tasks (Jin et al., 2024b; Luo et al., 2024b; Sun et al., 2024b; Ma et al., 2024; Jiang et al., 2024a; Yasunaga et al., 2021; Taunk et al., 2023).
@@ 5.2.3. Vector Indexing
5.2.4. Hybrid Indexing
Each of the above three indexing methods offers distinct advantages: graph indexing facilitates easy access to structural information, text indexing simplifies retrieval of textual content, and vector indexing enables quick and efficient searches. Therefore, in practical applications, a hybrid approach combining these indexing methods is often preferred over relying solely on one. For instance, HybridRAG (Sarmah et al., 2024) retrieves both vector and graph data simultaneously to enhance the content retrieved. While EWEK-QA (Dehghan et al., 2024) uses both text and knowledge graphs.
@@ 6.1.3. GNN-based Retriever
GNNs are adept at understanding and leveraging complex graph structures. GNN-based retrievers typically encode graph data and subsequently score different retrieval granularities based on their similarity to the query. For example, GNN-RAG (Mavromatis and Karypis, 2024) first encodes the graph, assigns a score to each entity, and retrieves entities relevant to the query based on a threshold. EtD (Liu et al., 2024e) iterates multiple times to retrieve relevant paths. During each iteration, it first uses LLaMA2 (Touvron et al., 2023) to select edges connecting the current node, then employs GNNs to obtain embeddings of the new layer of nodes for the next round of LLM selection.
6.1.4. Discussion
During the retrieval process, non-parametric retrievers exhibit good retrieval efficiency, but they may suffer from inaccurate retrieval due to a lack of training on downstream tasks. Meanwhile, although LM-based retrievers and GNN-based retrievers offer higher retrieval accuracy, they require significant computational overhead. Considering this complementarity, many methods propose hybrid retrieval approaches to improve both retrieval efficiency and accuracy. Many approaches adopt a multi-stage retrieval strategy, employing different models at each stage. For example, RoG (Luo et al., 2024b) first utilizes LLMs to generate planning paths and then extracts paths satisfying the planning paths from knowledge graphs. GenTKGQA (Gao et al., 2024a) infers crucial relations and constraints from the query using LLMs and extracts triplets according to these constraints.
@@ 6.2.2. Iterative Retrieval
One distinctive characteristic of adaptive retrieval is to let models autonomously determine the optimal moments to finish the retrieval activities. For instance, (Wu et al., 2023b; Guo et al., 2024) leverage an LM for hop prediction, which serves as an indicator to end the retrieval. There is also a group of researchers who utilize model-generated special tokens or texts as termination signals for the retrieval process. For example, ToG (Sun et al., 2024b; Ma et al., 2024) prompts the LLM agent to explore the multiple possible reasoning paths until the LLM determines the question can be answered based on the current reasoning path. (Zhang et al., 2022b) trains a RoBERTa to expand a path from each topic entity. In the process, a virtual relation named as “[END]” is introduced to terminate the retrieval process.
@@ 6.2.3. Multi-Stage Retrieval
6.2.4. Discussion
In GraphRAG, once retrieval typically exhibits lower complexity and shorter response times, making it suitable for scenarios requiring real-time responsiveness. In contrast, iterative retrieval often involves higher time complexity, especially when employing LLMs as retrievers, potentially leading to longer processing times. However, this approach can yield higher retrieval accuracy by iteratively refining retrieved information and generating responses. Therefore, the choice of retrieval paradigm should balance accuracy and time complexity based on specific use cases and requirements.
@@ 6.3.1. Nodes
Nodes allow for precise retrieval focused on individual elements within the graph, which is ideal for targeted queries and specific information extraction. In general, for knowledge graphs, nodes refer to entities. For other types of text attribute graphs, nodes may include textual information that describes the node’s attributes. By retrieving nodes within the graph, GraphRAG systems could provide detailed insights into their attributes, relationships, and contextual information. For example,  Munikoti et al. (2023) , Li et al. (2024b) and Wang et al. (2024b) construct document graphs and retrieves relevant passage nodes.  Liu et al. (2024e) , Sun et al. (2019) and Gutiérrez et al. (2024) retrieve entities from constructed knowledge graphs.
@@ 6.3.3. Paths
However, path retrieval can be challenging due to the exponential growth in possible paths as graph size increases, which escalates computational complexity. To address this, some methods retrieve relevant paths based on pre-defined rules. For example,  Wang et al. (2024a) and Lo and Lim (2023) first select entity pairs in the query and then traverse to find all the paths between them within $n$-hop. HyKGE (Jiang et al., 2024a) first defines three types of paths: path, co-ancestor chain, and co-occurrence chain, and then utilizes corresponding rules to retrieve each of these three types of paths. In addition, some methods utilize models to perform path searching on graphs. ToG (Sun et al., 2024b; Ma et al., 2024) proposes to prompt the LLM agent to perform the beam search on KGs and find multiple possible reasoning paths that help answer the question.  Luo et al. (2024b) , Wu et al. (2023b) and Guo et al. (2024) first utilizes the model to generate faithful reasoning plans and then retrieves relevant paths based on these plans. GNN-RAG (Mavromatis and Karypis, 2024) first identifies the entities in the question. Subsequently, all paths between entities that satisfy a certain length relationship are extracted.
@@ 6.3.5. Hybrid Granularties
6.3.6. Discussion
(1) In real applications, there are no clear boundaries between these retrieval granularities, as subgraphs can be composed of multiple paths, and paths can be formed by several triplets. (2) Various granularities such as nodes, triplets, paths, and subgraphs offer distinct advantages in the GraphRAG process. Balancing between retrieval content and efficiency is crucial when selecting the granularity, depending on the specific context of the task. For straightforward queries or when efficiency is paramount, finer granularities such as entities or triplets may be preferred to optimize retrieval speed and relevance. In contrast, complex scenarios often benefit from a hybrid approach that combines multiple granularities. This approach ensures a more comprehensive understanding of the graph structure and relationships, enhancing the depth and accuracy of the generated responses. Thus, GraphRAG’s flexibility in granularity selection allows it to adapt effectively to diverse information retrieval needs across various domains.
@@ 6.4.1. Query Enhancement
Due to the generally short length of queries and their limited information content, query expansion aims to improve search results by supplementing or refining the original query with additional relevant terms or concepts.  Luo et al. (2024b) generate relation paths grounded by KGs with LLMs to enhance the retrieval query.  Cheng et al. (2024) adopt SPARQL to get all the aliases of the query entities from Wikidata to augment the retrieval queries, which capture lexical variations of the same entity.  Huang et al. (2023a) propose a consensus-view knowledge retrieval method to improve retrieval accuracy, which first discover semantically relevant queries, and then re-weight the original query terms to enhance the retrieval performance. HyKGE (Jiang et al., 2024a) utilizes a large model to generate the hypothesis output of the question, concatenating the hypothesis output with the query as input to the retriever. Golden-Retriever (An et al., 2024) first recognizes the jargon in the query and then retrieves explanations of the jargon as a supplement to the query.